OpenStudy (anonymous):

What are the steps to solving this problem? 0.000027 divided by 0.000009 I will fan and give medal if you help! I also need to know the exponants, thanks.

OpenStudy (tkhunny):

It may be simplest to see in Scientific Notation. \(\dfrac{0.000027}{0.000009} = \dfrac{27E-6}{9E-6} = 3E-6 = 0.000003\)

OpenStudy (anonymous):

ok thanks